OpenStudy (anonymous):

what is the inverse of y= 1/2x -2?

OpenStudy (abdullahm):

You just posted another question??

OpenStudy (abdullahm):

First, switch the x and y. So since we have y=1/2x - 2 We make it: x= 1/2 y -2 Now solve for y

OpenStudy (abdullahm):

Some sites first tell you to solve for x in your original equation and then switch the x or y. But it gives you the same answer nonetheless.

OpenStudy (anonymous):

wait so it would be y= 2x+4???

OpenStudy (anonymous):

oh sweet thanks

OpenStudy (anonymous):

btw I posted another question, because no one would answer it

OpenStudy (abdullahm):

You got it correct :)

OpenStudy (anonymous):

oh sweet thanks